For each chemical below, complete these steps: a. Identify the type of chemical (strong acid, weak acid, strong base, weak base, salt, or water) b. Write the chemical reaction that would take place with water and include the K value. c. Complete an ICE chart. d. Solve for the pH. Circle your answer. 1. 0.20 M NH20H 2. 0.45 M H 3. 0.04 M Li20
Answer #1

NH2OH= weak base

NH2OH+H2O=[NH3OH]^(+) + OH^(-)

Ka Value= 1×10^(-6)

At equilibrium

(0.2-x) for hydroxyl amine x for protonated hydroxyl amine + x for OH-

[H+]=?ka×C = ?1×10^(-6)× 0.2= 4.4×10^-4

pH=-log[H+]= -log(4.4× 10^(-4))= 3.35

HI + H2O = H3O+. + OH-

at equilibrium

0.45-x for HI. X for H3O+ and OH-

Ka= 10^(10)

H+ = ?ka×C = ? 10^(10)×0.45 = 6.7× 10^(-6)

Therefore pH = -log ( 6.7× 10^(-6))= 5.17.